Jordan could row at 3 times the speed of the water in the river. He could row 16 miles downstream in 2 hours less than it took him to row 14 miles upstream, what was Jordan's rowing speed in still water?
Answer by richwmiller(17219)

s*t=d
r=3w
(r+w)*(t-2)=16
(r-w)*t=14
Jordan's speed in still water
r=4.5 mph
extra info
current
w=1.5 mph
time
t=14/3 hrs. upstream
t=4 2/3 hrs upstream
2 2/3 hrs downstream